OpenStudy (fernandez555):

(01.02) Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21. −45 0 2 63

OpenStudy (shamil98):

Let f(x) =y We need to find y^-1( the inverse) To do so we switch the values of and y around and solve for y. \[\Huge x = 6(y+2)-3\] \[\Huge 21= 6(y+2)-3\] Let me know what you get for y.

OpenStudy (anonymous):

need to solve for y.